Skip to end of metadata
Go to start of metadata




  • Meeting recorder (Last few times were  Orion, Ken, Jen, Colin, Hsin-Fang, Colin): 
  • Announcements
  • Action Items last time
  •  Campaigns
    • Monthly RC2/DC2
      •  Orion sent ahead: My updates are basically that HSC-RC2 w19 ran in CM fully and metrics are on Sasquatch Chronograf ! Need to do the same for DC2 w21
    • HSC PDR2
      • Made configuration for WIDE step2a:  quantum graph generation requires a lot of resources (32GB of RAM and 4 hours). Have a decision to make: Ask for more memory or make the groups smaller. Unclear which one takes less human time. Probably asking for more memory.
      • Will make a PDR2/DEEP + PDR2/WIDE + PDR2/UDEEP = PDR2/ALL campaign. 
      • PDR2/ALL requires specific configuration details on confluence and JIRA
      • BY: On the "long tail problem:" Suggestion to look at how the pilots in PanDA connect to the workers.  Maybe they are timing out. Would it be useful if we monitor the worker nodes?   Work is ongoing to debut pilot / worker node interactions.   Production can continue, but please report long tails to #ops-cm-team or #ops-usdf.
    • LATISS 10am DRP
      • Recall that each campaign nominally has a little steering committee. Was asked about setting one up for this campaign. 
      • accumulates the past 2 runs. PREOPS-3463 - Getting issue details... STATUS  ticket. Included DIA steps in DM-39038 - Getting issue details... STATUS . Meredith is looking at the outputs. 
      • Formalize "committee & primary stakeholder"
    • LATISS Prompt Processing at the USDF
      • Success, modulo weather
      • New templates coming, subject to data availability
    • Monthly AP on HSC
      • Increased QC generation, maybe 2 / 3X.
      • New templates seem to be working well.
      • New error to investigate.
    • Calibration Production
      • Previous calibration done at the end of the month with the noRG production
      • Working on make new production.
  • Tooling:
  • Show and tell 
  • AOB - none